Suzan is 7years older than her brother ray.the combined total of their age is 41 hol is ray?
Answer by Maths68(1474)

let
Present age of Suzan = s
Present age of Ray = r

Suzan is 7years older than her brother Ray
s=r+7............(1)
Sum of their ages is 41
s+r=41.........(2)

Put the value of s from (1) to (2)

r+7+r=41
2r=41-7
2r=34
2r/2=34/2
r=17

Put the value of r in (1)
s=r+7
s=17+7
s=24

Present age of Suzan = s = 24 years old
Present age of Ray = r = 17 years old